OpenStudy (anonymous):

Determine two compatible numbers to use to estimate the quotient 5,071(division)62

OpenStudy (compassionate):

\[\frac{ 5,071 }{ 62 }\] Are you trying to find out what the LCM is of these two?

OpenStudy (anonymous):

That's why I'm confused

OpenStudy (anonymous):

Awnsers are 4,800 and 60 4,900 and 70 5,000 and 60

OpenStudy (compassionate):

I think a good estimate would be within the closest 100's and 10's of the units, so 62 is 60 if we round it, and 5,071 becomes 5,000 when rounded. It's more logical. Think of it this way: if I was dividing \[\frac{ 4 }{ 2 }\] a better estimate would be 5/3 instead of 7/4 because 5/3 is closer to 2/4 than 7/4

hero (hero):

Hint:\[6 \times 8 = 48\]